An automobile is pushed 120 inches by a woman exerting 80 lbs of force horizontally on the vehicle. How much work does she do (a) in ft ⋅ lb and (b) in joules?


1
Expert's answer
2022-05-18T15:27:59-0400
"W=Fd\cos\theta\\
=\rm 80\: lbs\times 10\: ft\times \cos0^\circ=800\: ft\cdot lb=1088\: J"